# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v2

@@ Abstract
The Probably Approximately Correct (PAC) Privacy framework [46] provides a powerful instance-based methodology to preserve privacy in complex data-driven systems. Existing PAC Privacy algorithms (we call them Auto-PAC) rely on a Gaussian mutual information upper bound. However, we show that the upper bound obtained by these algorithms is tight if and only if the perturbed mechanism output is jointly Gaussian with independent Gaussian noise. We propose two approaches for addressing this issue. First, we introduce two tractable post-processing methods for Auto-PAC, based on Donsker–Varadhan representation and sliced Wasserstein distances. However, the result still leaves "wasted" privacy budget. To address this issue more fundamentally, we introduce Residual-PAC (R-PAC) Privacy, an $f$-divergence-based measure to quantify privacy that remains after adversarial inference. To implement R-PAC Privacy in practice, we propose a Stackelberg Residual-PAC (SR-PAC) privatization mechanism, a game-theoretic framework that selects optimal noise distributions through convex bilevel optimization. Our approach achieves efficient privacy budget utilization for arbitrary data distributions and naturally composes when multiple mechanisms access the dataset. Through extensive experiments, we demonstrate that SR-PAC obtains consistently a better privacy-utility tradeoff than both PAC and differential privacy baselines.
@@ 1 Introduction
Machine-learning models power critical applications—from medical diagnosis to autonomous vehicles—yet their outputs can inadvertently expose sensitive training data. As pipelines grow in scale and complexity, practitioners need rigorous, scalable privacy guarantees that go beyond ad-hoc testing. Over the past two decades, formal privacy frameworks have proliferated. Differential Privacy (DP) [13] (and its variants such as Rényi DP [33]) delivers input-independent worst-case indistinguishability by bounding output shifts from single-record changes. Alternative information-theoretic definitions, such as mutual-information DP [10], Fisher-information bounds [16, 23, 21], and Maximal Leakage [26, 39], provide complementary guarantees and offer alternative trade-offs between privacy and utility.
Nevertheless, provable privacy guarantees for modern data-processing algorithms remains a challenge. First, worst-case frameworks like DP require computing global sensitivity, which is generally NP-hard [47]. Moreover, computing the optimal privacy bound of DP under composition is, in general, a #P-complete [34]. In practice, finding the minimal noise needed to meet a target guarantee is intractable for most real-world algorithms, especially when the effect of each operation on privacy is unclear. On the other hand, empirical or simulation-based methods (e.g., testing resistance to membership inference [42]) address specific threats but lack rigorous, adversary-agnostic assurance. Bridging this gap requires a new, broadly applicable framework that can quantify and enforce privacy risk without relying on sensitivity.
A promising alternative has recently emerged: the Probably Approximately Correct (PAC) Privacy framework [46]. PAC Privacy shifts from indistinguishability-based guarantees to an operational notion that measures the information-theoretic hardness of reconstructing sensitive data. It is defined by an impossibility-of-inference guarantee for a chosen adversarial task and data prior, and the framework provides algorithms that enforce tractable mutual-information upper bounds to certify this guarantee. This approach enables automatic privatization via black-box simulation, and enjoys additive composition bounds and automatic privacy budget implementations for adaptive sequential compositions of mechanisms with arbitrary interdependencies. Notably, PAC Privacy often requires only $O(1)$ noise magnitude to achieve its privacy guarantees—independent of the output dimension—whereas differential privacy’s worst-case, input-independent noise magnitude scales as $\Theta(\sqrt{d})$ for a $d$-dimensional release.
@@
We address this limitation of Auto-PAC in two ways. First, working within the general PAC Privacy framework, we develop two tractable post-processing methods for Auto-PAC, based on Donsker–Varadhan representation and sliced Wasserstein distances. However, even these methods fail to fully close the privacy budget gap. To address this issue more fundamentally, we introduce the notion of Residual-PAC Privacy (R-PAC privacy). Unlike PAC privacy, which aims to directly bound mutual information, R-PAC privacy focuses instead on quantifying privacy budget remaining after information has been leaked by a data processing mechanism, using $f$-divergence to this end. When $f$-divergence is instantiated as Kullback–Leibler (KL) divergence, we show that Residual-PAC Privacy is fully characterized by the conditional entropy up to a known constant that does not depend on the mechanism or the applied noise.
@@
  • We characterize the conservativeness of Auto-PAC [46, 43], showing that it arises from the gap between the Gaussian surrogate bound and the true non-Gaussian mutual information of the privatized mechanism.
@@ 1.1 Related Work
Privacy Quantification Notions. Quantitative notions of privacy leakage have been extensively studied across a variety of contexts, leading to mathematically rigorous frameworks for assessing the amount of sensitive information that can be inferred by adversaries. Differential privacy (DP) and its variants have become the gold standard for formal privacy guarantees, with the original definitions by Dwork et al. [14, 13] formalizing privacy loss through bounds on the distinguishability of outputs under neighboring datasets. Variants such as concentrated differential privacy (CDP) [5, 15], zero-concentrated DP (zCDP) [4], and Rényi differential privacy (RDP) [33] have further extended this framework by parameterizing privacy loss with different statistical divergences (e.g., R’enyi divergence), thereby enhancing flexibility in privacy accounting, especially for compositions and adaptive mechanisms. Information-theoretic measures provide alternative and complementary approaches for quantifying privacy loss. For instance, mutual information has been used to analyze privacy leakage in a variety of settings [7, 10], with $f$-divergence and Fisher information offering finer-grained or context-specific metrics [46, 16, 23, 21]. These frameworks help to bridge the gap between statistical risk and adversarial inference, and are closely connected to privacy-utility trade-offs in mechanism design. Maximal leakage, hypothesis testing privacy, and other relaxations further broaden the analytic toolkit for measuring privacy risk.
@@
Privacy-Utility Trade-off. Balancing the trade-off between privacy and utility is a central challenge in the design of privacy-preserving mechanisms. This challenge is frequently formulated as an optimization problem [30, 40, 32, 19, 22, 18, 12, 1, 20]. For example, Ghosh et al. [19] demonstrated that the geometric mechanism is universally optimal for differential privacy under certain loss-minimizing criteria in Bayesian settings, while Lebanon et al. [30] and Alghamdi et al. [1] studied utility-constrained optimization. Gupte et al. [22] modeled the privacy-utility trade-off as a zero-sum game between privacy mechanism designers and adversaries, illustrating the interplay between optimal privacy protection and worst-case loss minimization.
@@
Optimization Approaches for Privacy. A growing body of work frames the design of privacy-preserving mechanisms as explicit optimization problems, aiming to maximize data utility subject to formal privacy constraints. Many adversarial or game-theoretic approaches—such as generative adversarial privacy (GAP)[24] and related GAN-based frameworks[8, 35, 27]—cast the privacy mechanism designer and the adversary as players in a min-max game, optimizing utility loss and privacy leakage, respectively. More recently, Selvi et al. [41] introduced a rigorous optimization framework for differential privacy based on distributionally robust optimization (DRO), formulating the mechanism design problem as an infinite-dimensional DRO to derive noise-adding mechanisms that are nonasymptotically and unconditionally optimal for a given privacy level. Their approach yields implementable mechanisms via tractable finite-dimensional relaxations, often outperforming classical Laplace or Gaussian mechanisms on benchmark tasks. Collectively, these lines of research illustrate the power of optimization and game-theoretic perspectives in achieving privacy-utility trade-offs beyond conventional mechanism design.
@@ 2.1 PAC Privacy
The PAC privacy framework [46] addresses this threat model and is formally defined as follows.
Definition 1 ($(\delta,\rho,\mathcal{D})$-PAC Privacy [46]).
@@
This privacy definition is highly flexible by enabling $\rho$ to encode a wide range of threat models and user-specified risk criteria. For example, in membership inference attacks [6], $\rho(\tilde{X},X)=1$ may indicate that $\widetilde{X}$ successfully determines the presence of a target data point $u_{0}$ in $X$. In reconstruction attacks [2], success may be defined by $\rho(\widetilde{X},X)=1$ if $|\widetilde{X}-X|_{2}\leq 1$, representing a close approximation of the original data.
@@
Definition 3 ($(\Delta_{f}^{\delta},\rho,\mathcal{D})$ PAC Advantage Privacy [46]).
@@ 2.2 Automatic PAC Privatization Algorithms
PAC Privacy enables automatic privatization, which supports simulation-based implementation for arbitrary black-box mechanisms, without requiring the worst-case adversarial analysis, such as sensitivity computation. In this section, we present the main theorems and algorithms underlying automatic PAC privatization as introduced in [46] (hereafter "Auto-PAC") and the efficiency-improved version proposed in [43] (hereafter "Efficient-PAC"; algorithm details in Appendix A). We start by defining the mutual information.
@@
When the $f$-divergence in $\Delta_{f}^{\delta}$ is instantiated as the KL divergence (denoted as $\Delta_{\mathrm{KL}}^{\delta}$), Theorem 1 of [46] shows
@@
Algorithm 1 $(1-\gamma)$-Confidence Auto-PAC [46]
@@
Theorem 1 establishes a simple upper bound on the mutual information with Gaussian noise perturbation. Choosing $\Sigma_{B}$ to implement the Gaussian surrogate bound $\mathtt{LogDet}(\mathcal{M}(X),B)=\beta$ for a privacy budget $\beta$ enables anisotropic noise as it estimates the eigenvectors of $\mathcal{M}(X)$ to fit the instance-based noise to the geometry of the eigenspace of $\mathcal{M}(X)$. The result extends naturally to randomized mechanisms of the form $\mathcal{M}(X,\theta)$, where $\theta$ is a random seed (Corollary 2 of [46]). Building on Theorem 1, an automatic PAC privatization algorithm (Auto-PAC) shown in Algorithm 1 is proposed by [46] to determine an appropriate Gaussian noise covariance $\Sigma_{B}$ to ensure that $\mathtt{MI}(X;\mathcal{M}(X)+B)\leq\beta$ with confidence at least $1-\gamma$. This is achieved using the user-specific security parameter $c$, privacy budget partitions $v$ and $\beta^{\prime}$ such that $\beta=v+\beta^{\prime}$ (Theorem 4 of [46]). We refer to Algorithm 1 as $(1-\gamma)$-Confidence Auto-PAC.
@@ 2.3 Differential Privacy
Definition 5 ($(\epsilon,\bar{\delta})$-Differential Privacy [14]).
@@ 3.1 Mechanism Comparison in PAC Privacy
Definition 9 of [46] defines the optimal perturbation that tightly implements the privacy budget while maintaining optimal utility, where utility is captured by a loss function $\mathcal{K}$. An optimal perturbation $Q^{*}$ is a solution of the following optimization problem:
@@ 3.2 $\mathtt{Gap}_{\mathtt{d}}$ Reduction via Non-Gaussianity Correction
Definition 6 (Donsker–Varadhan (DV) Objective [11]).
@@
where the supremum ranges over measurable $f$ such that $\mathbb{E}_{Q}[e^{f}]<\infty$. We call $\mathcal{J}(f;P,Q)\equiv\mathbb{E}_{P}[f]-\log\mathbb{E}_{Q}[e^{f}]$ the DV objective. In our setting, $\mathtt{D}_{Z}=\mathtt{D}_{\mathrm{KL}}(P_{Z}\|\widetilde{Q}_{\mathcal{M}})=\sup_{f}\mathcal{J}\big{(}f;P_{\mathcal{M},B},\widetilde{Q}_{\mathcal{M}}\big{)}$.
@@
$\widehat{\mathtt{D}}_{\mathrm{LCE}}\equiv\Big{[}\widehat{\mathcal{J}}\big{(}\widehat{f}_{\mathrm{tr}};S_{P}^{\mathrm{val}},S_{Q}^{\mathrm{val}}\big{)}-\Gamma_{\hat{\delta}}\Big{]}_{+}$.
@@
    $0\leq\sup_{f}\widehat{\mathtt{D}}_{Z}(f)=\mathtt{D}_{\mathrm{KL}}\big{(}P_{\mathcal{M},B}\big{\|}\widetilde{Q}_{\mathcal{M}}\big{)}$.
@@
$\widehat{\mathtt{D}}_{Z}\equiv\frac{1}{2\lambda_{\max}(\Sigma_{Z})}\mathrm{SW}_{2}^{2}\big{(}P_{\mathcal{M},B},\widetilde{Q}_{\mathcal{M}}\big{)}$.

### v3

@@ Abstract
The Probably Approximately Correct (PAC) Privacy framework [xiao2023pac] provides a powerful instance-based methodology to preserve privacy in complex data-driven systems. Existing PAC Privacy algorithms (we call them Auto-PAC) rely on a Gaussian mutual information upper bound. However, we show that the upper bound obtained by these algorithms is tight if and only if the perturbed mechanism output is jointly Gaussian with independent Gaussian noise. We propose two approaches for addressing this issue. First, we introduce two tractable post‐processing methods for Auto-PAC, based on Donsker–Varadhan representation and sliced Wasserstein distances. However, the result still leaves "wasted" privacy budget. To address this issue more fundamentally, we introduce Residual-PAC (R-PAC) Privacy, an $f$-divergence-based measure to quantify privacy that remains after adversarial inference. To implement R-PAC Privacy in practice, we propose a Stackelberg Residual-PAC (SR-PAC) privatization mechanism, a game-theoretic framework that selects optimal noise distributions through convex bilevel optimization. Our approach achieves efficient privacy budget utilization for arbitrary data distributions and naturally composes when multiple mechanisms access the dataset. Through extensive experiments, we demonstrate that SR-PAC obtains consistently a better privacy-utility tradeoff than both PAC and differential privacy baselines.
@@ 1 Introduction
Machine‐learning models power critical applications—from medical diagnosis to autonomous vehicles—yet their outputs can inadvertently expose sensitive training data. As pipelines grow in scale and complexity, practitioners need rigorous, scalable privacy guarantees that go beyond ad-hoc testing. Over the past two decades, formal privacy frameworks have proliferated. Differential Privacy (DP) [dwork2006differential] (and its variants such as Rényi DP [mironov2017renyi]) delivers input-independent worst-case indistinguishability by bounding output shifts from single-record changes. Alternative information-theoretic definitions, such as mutual-information DP [cuff2016differential], Fisher-information bounds [farokhi2017fisher, hannun2021measuring, guo2022bounding], and Maximal Leakage [issa2019operational, saeidian2023pointwise], provide complementary guarantees and offer alternative trade-offs between privacy and utility.
Nevertheless, provable privacy guarantees for modern data‐processing algorithms remains a challenge. First, worst‐case frameworks like DP require computing global sensitivity, which is generally NP-hard [xiao2008output]. Moreover, computing the optimal privacy bound of DP under composition is, in general, a #P-complete [murtagh2015complexity]. In practice, finding the minimal noise needed to meet a target guarantee is intractable for most real-world algorithms, especially when the effect of each operation on privacy is unclear. On the other hand, empirical or simulation-based methods (e.g., testing resistance to membership inference [shokri2017membership]) address specific threats but lack rigorous, adversary-agnostic assurance. Bridging this gap requires a new, broadly applicable framework that can quantify and enforce privacy risk without relying on sensitivity.
A promising alternative has recently emerged: the Probably Approximately Correct (PAC) Privacy framework [xiao2023pac]. PAC Privacy shifts from indistinguishability-based guarantees to an operational notion that measures the information-theoretic hardness of reconstructing sensitive data. It is defined by an impossibility-of-inference guarantee for a chosen adversarial task and data prior, and the framework provides algorithms that enforce tractable mutual-information upper bounds to certify this guarantee. This approach enables automatic privatization via black-box simulation, and enjoys additive composition bounds and automatic privacy budget implementations for adaptive sequential compositions of mechanisms with arbitrary interdependencies. Notably, PAC Privacy often requires only $O(1)$ noise magnitude to achieve its privacy guarantees—independent of the output dimension—whereas differential privacy’s worst‐case, input‐independent noise magnitude scales as $\Theta(\sqrt{d})$ for a $d$-dimensional release.
@@
We address this limitation of Auto-PAC in two ways. First, working within the general PAC Privacy framework, we develop two tractable post‐processing methods for Auto-PAC, based on Donsker–Varadhan representation and sliced Wasserstein distances. However, even these methods fail to fully close the privacy budget gap. To address this issue more fundamentally, we introduce the notion of Residual-PAC Privacy (R-PAC privacy). Unlike PAC privacy, which aims to directly bound mutual information, R-PAC privacy focuses instead on quantifying privacy budget remaining after information has been leaked by a data processing mechanism, using $f$-divergence to this end. When $f$-divergence is instantiated as Kullback–Leibler (KL) divergence, we show that Residual-PAC Privacy is fully characterized by the conditional entropy up to a known constant that does not depend on the mechanism or the applied noise.
@@
  • We characterize the conservativeness of Auto-PAC [xiao2023pac, sridhar2024pac], showing that it arises from the gap between the Gaussian surrogate bound and the true non-Gaussian mutual information of the privatized mechanism.
@@ 1.1 Related Work
Privacy Quantification Notions. Quantitative notions of privacy leakage have been extensively studied across a variety of contexts, leading to mathematically rigorous frameworks for assessing the amount of sensitive information that can be inferred by adversaries. Differential privacy (DP) and its variants have become the gold standard for formal privacy guarantees, with the original definitions by Dwork et al. [dwork2006calibrating, dwork2006differential] formalizing privacy loss through bounds on the distinguishability of outputs under neighboring datasets. Variants such as concentrated differential privacy (CDP) [bun2016concentrated, dwork2016concentrated], zero-concentrated DP (zCDP) [bun2018composable], and Rényi differential privacy (RDP) [mironov2017renyi] have further extended this framework by parameterizing privacy loss with different statistical divergences (e.g., R’enyi divergence), thereby enhancing flexibility in privacy accounting, especially for compositions and adaptive mechanisms. Information-theoretic measures provide alternative and complementary approaches for quantifying privacy loss. For instance, mutual information has been used to analyze privacy leakage in a variety of settings [chatzikokolakis2010statistical, cuff2016differential], with $f$-divergence and Fisher information offering finer-grained or context-specific metrics [xiao2023pac, farokhi2017fisher, hannun2021measuring, guo2022bounding]. These frameworks help to bridge the gap between statistical risk and adversarial inference, and are closely connected to privacy-utility trade-offs in mechanism design. Maximal leakage, hypothesis testing privacy, and other relaxations further broaden the analytic toolkit for measuring privacy risk.
@@
Privacy-Utility Trade-off. Balancing the trade-off between privacy and utility is a central challenge in the design of privacy-preserving mechanisms. This challenge is frequently formulated as an optimization problem [lebanon2009beyond, sankar2013utility, lopuhaa2024mechanisms, ghosh2009universally, gupte2010universally, geng2020tight, du2012privacy, alghamdi2022cactus, goseling2022robust]. For example, Ghosh et al. [ghosh2009universally] demonstrated that the geometric mechanism is universally optimal for differential privacy under certain loss-minimizing criteria in Bayesian settings, while Lebanon et al. [lebanon2009beyond] and Alghamdi et al. [alghamdi2022cactus] studied utility-constrained optimization. Gupte et al. [gupte2010universally] modeled the privacy-utility trade-off as a zero-sum game between privacy mechanism designers and adversaries, illustrating the interplay between optimal privacy protection and worst-case loss minimization.
@@
Optimization Approaches for Privacy. A growing body of work frames the design of privacy-preserving mechanisms as explicit optimization problems, aiming to maximize data utility subject to formal privacy constraints. Many adversarial or game-theoretic approaches—such as generative adversarial privacy (GAP)[huang2018generative] and related GAN-based frameworks[chen2018understanding, nasr2018machine, jordon2018pate]—cast the privacy mechanism designer and the adversary as players in a min-max game, optimizing utility loss and privacy leakage, respectively. More recently, Selvi et al. [selvi2025differential] introduced a rigorous optimization framework for differential privacy based on distributionally robust optimization (DRO), formulating the mechanism design problem as an infinite-dimensional DRO to derive noise-adding mechanisms that are nonasymptotically and unconditionally optimal for a given privacy level. Their approach yields implementable mechanisms via tractable finite-dimensional relaxations, often outperforming classical Laplace or Gaussian mechanisms on benchmark tasks. Collectively, these lines of research illustrate the power of optimization and game-theoretic perspectives in achieving privacy-utility trade-offs beyond conventional mechanism design.
@@ 2.1 PAC Privacy
The PAC privacy framework [xiao2023pac] addresses this threat model and is formally defined as follows.
Definition 1 ($(\delta,\rho,\mathcal{D})$-PAC Privacy [xiao2023pac]).
@@
This privacy definition is highly flexible by enabling $\rho$ to encode a wide range of threat models and user-specified risk criteria. For example, in membership inference attacks [carlini2022membership], $\rho(\tilde{X},X)=1$ may indicate that $\widetilde{X}$ successfully determines the presence of a target data point $u_{0}$ in $X$. In reconstruction attacks [balle2022reconstructing], success may be defined by $\rho(\widetilde{X},X)=1$ if $|\widetilde{X}-X|_{2}\leq 1$, representing a close approximation of the original data.
@@
Definition 3 ($(\Delta_{f}^{\delta},\rho,\mathcal{D})$ PAC Advantage Privacy [xiao2023pac]).
@@ 2.2 Automatic PAC Privatization Algorithms
PAC Privacy enables automatic privatization, which supports simulation-based implementation for arbitrary black-box mechanisms, without requiring the worst-case adversarial analysis, such as sensitivity computation. In this section, we present the main theorems and algorithms underlying automatic PAC privatization as introduced in [xiao2023pac] (hereafter "Auto-PAC") and the efficiency-improved version proposed in [sridhar2024pac] (hereafter "Efficient-PAC"; algorithm details in Appendix A). We start by defining the mutual information.
@@
When the $f$-divergence in $\Delta_{f}^{\delta}$ is instantiated as the KL divergence (denoted as $\Delta_{\mathrm{KL}}^{\delta}$), Theorem 1 of [xiao2023pac] shows
@@
Algorithm 1 $(1-\gamma)$-Confidence Auto-PAC [xiao2023pac]
@@
Theorem 1 establishes a simple upper bound on the mutual information with Gaussian noise perturbation. Choosing $\Sigma_{B}$ to implement the Gaussian surrogate bound $\mathtt{LogDet}(\mathcal{M}(X),B)=\beta$ for a privacy budget $\beta$ enables anisotropic noise as it estimates the eigenvectors of $\mathcal{M}(X)$ to fit the instance-based noise to the geometry of the eigenspace of $\mathcal{M}(X)$. The result extends naturally to randomized mechanisms of the form $\mathcal{M}(X,\theta)$, where $\theta$ is a random seed (Corollary 2 of [xiao2023pac]). Building on Theorem 1, an automatic PAC privatization algorithm (Auto-PAC) shown in Algorithm 1 is proposed by [xiao2023pac] to determine an appropriate Gaussian noise covariance $\Sigma_{B}$ to ensure that $\mathtt{MI}(X;\mathcal{M}(X)+B)\leq\beta$ with confidence at least $1-\gamma$. This is achieved using the user-specific security parameter $c$, privacy budget partitions $v$ and $\beta^{\prime}$ such that $\beta=v+\beta^{\prime}$ (Theorem 4 of [xiao2023pac]). We refer to Algorithm 1 as $(1-\gamma)$-Confidence Auto-PAC.
@@ 2.3 Differential Privacy
Definition 5 ($(\epsilon,\bar{\delta})$-Differential Privacy [dwork2006calibrating]).
@@ 3.1 Mechanism Comparison in PAC Privacy
Definition 9 of [xiao2023pac] defines the optimal perturbation that tightly implements the privacy budget while maintaining optimal utility, where utility is captured by a loss function $\mathcal{K}$. An optimal perturbation $Q^{*}$ is a solution of the following optimization problem:
@@ 3.2 $\mathtt{Gap}_{\mathtt{d}}$ Reduction via Non-Gaussianity Correction
Definition 6 (Donsker–Varadhan (DV) Objective [donsker1975asymptotic]).
@@
[…]
@@
$\widehat{\mathtt{D}}_{\mathrm{LCE}}\equiv\Big[\widehat{\mathcal{J}}\big(\widehat{f}_{\mathrm{tr}};S_{P}^{\mathrm{val}},S_{Q}^{\mathrm{val}}\big)-\Gamma_{\hat{\delta}}\Big$$_{+}$.
@@
[…]
@@
[…]